# Progressive Growing of Video Tokenizers for Highly Compressed Latent Spaces

### v1

@@ Abstract
Video tokenizers are essential for latent video diffusion models, converting raw video data into spatiotemporally compressed latent spaces for efficient training. However, extending state-of-the-art video tokenizers to achieve a temporal compression ratio beyond 4× without increasing channel capacity poses significant challenges. In this work, we propose an alternative approach to enhance temporal compression. We find that the reconstruction quality of temporally subsampled videos from a low-compression encoder surpasses that of high-compression encoders applied to original videos. This indicates that high-compression models can leverage representations from lower-compression models. Building on this insight, we develop a bootstrapped high-temporal-compression model that progressively trains high-compression blocks atop well-trained lower-compression models. Our method includes a cross-level feature-mixing module to retain information from the pretrained low-compression model and guide higher-compression blocks to capture the remaining details from the full video sequence. Evaluation of video benchmarks shows that our method significantly improves reconstruction quality while increasing temporal compression compared to direct extensions of existing video tokenizers. Furthermore, the resulting compact latent space effectively trains a video diffusion model for high-quality video generation with a reduced token budget.
Figure 1: (left) Our video tokenizer ProMAG is capable of much more effectively reconstructing high motion videos even at very large compression ratios, like 16× temporal compression compared to baseline. Dotted boxes highlight regions where the baseline fails catastrophically. (right) We show that our highly compressed latent space (16×) is suitable for training text-to-video diffusion models.
footnotetext: This work was done during an internship at Adobe Research.
@@ 1 Introduction
The emergence of diffusion models [15, 36] has transformed image [33, 30, 9, 32, 27] and video [10, 4, 14, 35, 16, 40, 5] generation. These models gradually transform random noise into coherent visual outputs, showcasing impressive capabilities in producing high-fidelity content [5, 8]. Latent diffusion models (LDMs)[32] have gained popularity for image and video generation[33, 1]. Unlike pixel diffusion models that operate directly on raw pixels, LDMs project pixels into a low-dimensional latent space using a variational autoencoder (VAE) [23], enabling diffusion in this reduced space. This dimensionality reduction enhances computational and memory efficiency, which is particularly crucial for video generation due to the higher dimensionality of video data compared to images.
Early latent video diffusion models (LVDMs) relied on latents extracted from each frame independently [35, 14, 2] by reusing the same VAE from image LDMs. This neglects the temporal dynamics inherent in video data, compromising the temporal consistency of the generated videos, and offers no temporal compression in the latent space. To address these limitations, the seminal work MagViT-v2 [41] pioneers a dedicated spatiotemporal video encoding that can jointly encode images and videos in the same latent space. Originally designed to encode video into discrete spatiotemporal tokens, MagViT-v2 has been widely adapted to continuous tokens for use in many state-of-the-art LVDMs [13, 40, 46, 24].
@@
Interestingly, we observe that 4$\times$ temporal compression MagViT-v2 can accurately reconstruct videos with 4$\times$ frame subsampling (i.e., low FPS), much better than a 16$\times$ temporal compression model on the original (high FPS) video. Thus, we try to answer the question: “Can we meaningfully boost the temporal compression of video tokenizers by reusing a pretrained 4$\times$ model, while keeping the number of latent channels constant?” In this paper, we intentionally keep the latent channel dimension fixed over the progressive growing process for two reasons. First, that helps separate the effects of channel dimensions and progressive growing on the model performance. Second, one of our main goals is to achieve good latent features not only for high reconstruction quality but also for effective downstream diffusion model training. It has been observed that increasing the latent channel dimensions, while leading to a slight improvement in reconstruction, tends to make generative model training more difficult [41].
@@
  • We present our progressive model growing technique to train ProMAG for high temporal compression ratios, achieving up to 8$\times$ and even 16$\times$. Notably, we are the first to achieve high-quality reconstruction with a 16$\times$ temporal compression model. Evaluations on video benchmarks show that our models deliver significantly better reconstruction quality than simply extending existing video tokenizers for high compression.
  • We showcase the effectiveness of our highly compressed latents for text-to-video generation with DiT. Experiments on the text-to-video evaluation benchmark [18] reveal that DiT trained with our 16$\times$ temporal-compression latents achieves comparable or slightly higher quality than the standard 4$\times$ compression, while significantly enhancing efficiency.
@@ 2 Related work
Latent video diffusion models. The pioneering LVDMs [2, 35, 14, 10, 3, 26] repurposed image VAEs to transform each frame of the video independently into the latent space. This caused temporal flickering in the generated video and limited the video length due to limited compression achieved by image VAEs. More recent works [13, 40, 46, 24, 5] use VAEs that operate on video-level, hence mitigating both limitations of using image VAEs for video diffusion models. Building upon this, we aim to create a VAE with even higher temporal compression (8$\times$ and 16$\times$) compared to the contemporary 4$\times$, enabling highly efficient and longer video generation.
@@
High compression VAEs. There have been few works recently that investigate very high compression in latent space for higher-resolution (2K or 4K) images and longer video generation. UltraPixel [31] provides an image autoencoder that can perform 24$\times$ spatial compression. Concurrent work DC-AE [7] provides a solution to increase the spatial compression to as large as 128$\times$ spatially. However, they have to also increase channel dimension to cope with degradation in reconstruction quality. In comparison, we only focus on increasing the temporal compression, while keeping the number of latent channels fixed to avoid making the latent space harder to learn for the diffusion model. PVDM [42] treats video as a 3D volume and encodes it into a tri-plane-like representation. CMD [43] and LaMD [17] disentangle video into content and motion components, enabling generation with image-based diffusion models. Although this enables relatively higher compression, it is not clear how to extend such factorized latent space design to fit contemporary spatiotemporal VDMs. Similarly, concurrent work MovieGen [28] achieves 8$\times$ temporal compression. In comparison, our method can achieve 16$\times$ temporal compression.
@@
Figure 2: Motivation. We highlight the key motivation of our progressive growing approach. Directly training MagViT-v2 for 16$\times$ temporal compression leads to poor reconstruction quality for a $24$-fps video (bottom). $s_{f}=1$ stands for frame subsampling factor = 1. However, we observed that the 4$\times$ temporal compression model can still accurately reconstruct a $6$-fps video by feeding the same $24$-fps video after subsampling frames by a factor of 4×4$\times$4 ×, $s_{f}=4$ (top). This observation implies that it is not necessarily the large motion that leads to worse reconstruction, but that training all the weights of a much larger number of encoder and decoder blocks makes training unstable. Since our base model ProMAG at 4$\times$ temporal compression is similar to MagViT-v2, this idea is conclusion applicable for ProMAG.
@@
Hierarchical models. Hierarchical generation has been explored most notably using generative adversarial networks (GANs) [12] to perform high-resolution image generation in stages [39, 20, 19]. These approaches first learn a generator that operates in low resolution, followed by learning additional high-resolution blocks on top in stages. Similar ideas have also been explored in image and video generation using diffusion models [31, 1, 16, 9] where content is first generated in low-resolution and then upsampled in the spatial and temporal domain using a separate model. In our work, we are inspired by ProGAN [19] and adapt the progressive learning idea to our problem of boosting the temporal compression in video tokenizers, progressively learning separate model blocks to handle different levels of compression.
@@ 3 Method
Our goal is to create a continuous-token video tokenizer with high temporal compression ratios (8$\times$ or 16$\times$) and high reconstruction quality. In Section 3.1, we discuss the technical modifications we make to the MagViT-v2 implementation [41] to arrive at our (ProMAG) base model. In Section 3.2, we analyze why the base ProMAG, capable of doing 4$\times$ temporal compression, has difficulties extending to 8$\times$ or 16$\times$ temporal compression. We discuss a method of progressively growing the ProMAG-4X model to achieve 8$\times$ and consequently 16$\times$ temporal compression in Section 3.2. Finally, in Section 3.3 we introduce our layer-wise spatial tiling technique during decoding to enable high-resolution encoding-decoding.
@@ 3.1 Base model: ProMAG
Efficient upsampling. MagViT-v2 encodes a 17-frame video into $5$ latent frames which are decoded to $20$ frames after 4$\times$ temporal upsampling. The presence of three additional frames in the reconstruction is a byproduct of the fact that the first latent frame only represents a single input frame. The MagViT-v2 decoder discards these $3$ reconstruction frames when generating the final output. We find this operation to be wasteful in terms of computation and becomes a greater problem with more temporal compression. For example, in the case of 16$\times$ compression on $17$ frames, the decoder will output $32$ frames and have to reject the first $15$ frames. To mitigate this issue, in all the temporal upsampling layers of the decoder, we discard the $1^{st}$ frame, as it can be considered a padding frame in causal convolutions. This reduces the memory consumption significantly, especially for higher compression models, without hurting reconstruction quality.
Removing ‘spot’ artifacts. MagViT-v2 uses group normalization to normalize activations between layers. This causes ‘spot’ like artifacts when trained to progressively increase the temporal compression (Figure 3(a)). This occurs because the model tends to encode important global information in these high-norm latent pixels, as also observed in [28]. To address this phenomenon, we modify the group-normalization operation by removing the mean subtraction following [21, 22]. We found this modification effectively resolves the ‘spot’ artifacts in the reconstructed videos (Figure 3(b)).
@@ 3.2 Progressive model growing
Motivation. To increase the temporal compression of ProMAG from 4$\times$ to 8$\times$ or 16$\times$, the standard approach is to add additional downsampling and upsampling blocks in the bottleneck layers of the encoder and decoder respectively. In our preliminary study, we found that directly training ProMAG for 16$\times$ temporal compression leads to poor reconstruction quality for a $24$-fps video (Figure 2(b)). However, we observed that the 4$\times$ temporal compression model can still accurately reconstruct a $6$-fps video by feeding the same $24$-fps video after subsampling frames by a factor of 4×4$\times$4 × (Figure 2(a)). This observation implies that it is not necessarily the large motion that leads to worse reconstruction, but that training all the weights of a much larger number of encoder and decoder blocks makes training unstable.
Since the ProMAG-4$\times$ can accurately reconstruct a $6$-fps video, our strategy is to use it as guidance for the additional bottleneck downsampling and upsampling blocks to achieve 8$\times$ and 16$\times$ temporal compression. This can be thought of as a guided video interpolation problem. Instead of forcing the bottleneck compression layers to learn the entire information flow, we aim to induce them to reuse the information of the 4$\times$ blocks, and only learn the essential information needed to synthesize the in-between frames, which is necessary to reconstruct the full video.
@@
Key-frame embedding. A naive way to achieve 8$\times$ temporal compression using the 4$\times$ model, $\mathbf{E}_{4\times}$, is to subsample the input video frames $\mathbf{v}$ by a factor of two, ($\mathbf{v}_{//2}$), and then encode the frames with $\mathbf{E}_{4\times}$. This can be referred to as learning the key-frame encodings $z_{\text{key}}$, of the subsampled keyframes.
@@
where $\mathbf{\hat{E}}_{4\times}$ is the encoder blocks without the bottleneck $1\times 1\times 1$ layers.
Residual embedding. Now with the encodings to accurately reconstruct keyframes in the video, we only need to learn the information for the remaining frames. We denote $\overset{*}{z}$ as the embedding of entire video $\mathbf{v}$ upto the 4$\times$ temporal compression blocks, $\mathbf{\hat{E}}_{4\times}$. To reach 8$\times$ compression, we compress $\overset{*}{z}$ through the additional 2$\times$ downsampling block. However, we need to ensure that the downsampling blocks $\mathbf{\hat{E}}_{4\times\text{-}8\times}$ are aware of the information already present in $z_{\text{key}}$, as it only needs to preserve the remaining information to avoid redundancy. For this, we use adaptive group normalization (AdaNorm) to condition the intermediate latent, $\overset{*}{z}$, on the key-frame embeddings $z_{\text{key}}$ before passing it to $\mathbf{\hat{E}}_{4\times\text{-}8\times}$.
@@
For the decoder we only add an additional bottleneck upsampling block.
@@
We initialize the weights for the 8$\times$ temporal compression model with the 4$\times$ models and freeze the encoder and decoder blocks initialized from the 4$\times$ model. At this stage, we only train the newly added blocks and the $1\times 1\times 1$ bottleneck layer using (1) the $L_{1}$ reconstruction loss $\mathcal{L}_{\text{rec}}$, and (2) the KL-divergence loss $\mathcal{L}_{\text{KL}}$. While GAN loss can slightly improve the FVD score in the reconstructed videos, it tends to introduce more visual artifacts in challenging scenarios. Therefore, we do not use it at this stage, which also makes our progressive growing process significantly more efficient.
@@ 3.3 High resolution video reconstruction
Directly using our model, ProMAG, or MagViT-v2 to encode and decode videos at high resolution (e.g. 540$\times$960), results in out-of-memory (OOM) errors due to the VRAM-intensive 3D convolution layers.
@@
Figure 6: Qualitative Comparison with Baseline. Video reconstruction example for (left) 16$\times$ temporal compression, (right) 8$\times$ temporal compression. In both compression ratios, baseline MagViT-v2 causes loss of details and produces artifacts in the reconstructed frames. The dotted boxes show regions of the most difference between our results and baseline.
Method Compression $z\_dim$ MCL-JCV DAVIS
512$\times$512 360$\times$640 512$\times$512 360$\times$640
PSNR LPIPS rFVD PSNR LPIPS rFVD PSNR LPIPS rFVD PSNR LPIPS rFVD
OS-VAE 8$\times$8$\times$4 4 30.59 11.72 120.93 29.92 12.9 168.44 27.73 15.99 405.28 26.58 17.22 424.78
OSP-VAE 30.21 11.72 120.93 29.57 8.19 122.18 27.29 10.74 298.02 26.08 11.19 327.64
[…]
CMD (p2) 23.50 17.36 1120.30 - - - 17.72 28.31 2136.59 - - -
[…]
ProMAG 30.99 6.53 57.91 30.37 6.53 58.14 28.43 7.38 130.56 27.07 7.67 152.07
MagViT-v2 8$\times$8$\times$8 8 28.47 11.17 219.36 28.33 10.73 222.09 24.02 16.7 525.1 23.82 15.55 477.19
ProMAG 30.26 8.36 96.85 29.84 8.29 108.95 26.48 11.37 360.54 25.76 11.12 337.15
MagViT-v2 8$\times$8$\times$16 8 26.15 14.54 255.51 25.96 14.19 265.99 21.82 20.54 728.04 21.73 19.43 695.67
ProMAG 28.31 11.2 183.4 28.27 10.74 206.3 23.89 16.9 608.43 23.71 15.74 545.87
ProMAG 8$\times$8$\times$4 16 32.94 4.91 34.21 32.3 4.84 39.61 30.77 4.98 93.92 29.11 5.35 83.03
MagViT-v2 8$\times$8$\times$8 16 28.51 10.62 135.57 29.92 7.16 84.1 27.18 8.84 243.19 26.08 9.07 260.4
ProMAG 32.06 6.94 63.64 31.61 6.27 63.59 28.7 8.32 194.79 27.55 8.39 214.47
ProMAG (w/ GAN) 31.78 6.21 53.18 31.29 6.04 59.9 28.21 7.76 160.22 27.08 7.75 197.18
MagViT-v2 8$\times$8$\times$16 16 28.51 10.62 135.57 28.21 10.26 141.6 24.17 14.84 427.52 23.65 14.39 393.83
ProMAG 30.02 9.24 115.25 29.83 8.72 118.17 25.64 13.89 320.8 25.21 12.99 304.21
ProMAG (w/ GAN) 29.67 8.38 99.64 29.43 7.99 99.63 25.05 12.04 274.24 24.53 11.33 289.24
Table 1: Quantitative Comparison with Baselines on Reconstruction Quality. Comparison with baseline methods at different compression rates and number of latent channels on MCL-JCV and DAVIS datasets. The best results are bold-faced. CMD cannot generalize to non-square aspect ratios due to its architecture. Our method achieves better reconstruction quality under all scenarios compared to baseline methods. Note: CMD with p8 and p2 are models trained with 2 different patch-sizes (according to the author’s suggestion and official repository).
We found the OOM issue arises only during decoding (i.e., high-resolution videos can be encoded as-is). With this in mind, we opted to encode the entire video at once and only spatially tile the encoded latent, decoding each tile separately. However, similar to tiling in RGB space, this approach results in artifacts, as seen in Figure 5 (a). In the latent tiling case, we found the issue to be a misalignment between the full, low-resolution video encoding and decoding regimen learned during training and the tiled decoding seen at inference. Specifically, we found that there was a meaningful difference in normalization statistics encountered during training and inference. To mitigate this discrepancy, we introduce a layer-wise spatial tiling technique: during decoding, we divide the input to each conv layer into overlapping spatial tiles, process each tile independently, and merge the tiles back together using linear interpolation weights to blend the overlapping components. Using this approach, we can decode into high-resolution videos without any artifacts (Figure 5 (b)).

### v2

# Progressive Growing of Video Tokenizers for Temporally Compact Latent Spaces

@@ Abstract
Video tokenizers are essential for latent video diffusion models, converting raw video data into spatiotemporally compressed latent spaces for efficient training. However, extending state-of-the-art video tokenizers to achieve a temporal compression ratio beyond 4× without increasing channel capacity poses significant challenges. In this work, we propose an alternative approach to enhance temporal compression. We find that the reconstruction quality of temporally subsampled videos from a low-compression encoder surpasses that of high-compression encoders applied to original videos. This indicates that high-compression models can leverage representations from lower-compression models. Building on this insight, we develop a bootstrapped high-temporal-compression model that progressively trains high-compression blocks atop well-trained lower-compression models. Our method includes a cross-level feature-mixing module to retain information from the pretrained low-compression model and guide higher-compression blocks to capture the remaining details from the full video sequence. Evaluation of video benchmarks shows that our method significantly improves reconstruction quality while increasing temporal compression compared to directly training the full model. Furthermore, the resulting compact latent space effectively trains a video diffusion model for high-quality video generation with a significantly reduced token budget.
Figure 1: (left) Our video tokenizer ProMAG is capable of much more effectively reconstructing high motion videos even at very large (16$\times$) temporal compression, compared to baseline MagViT-v2. Dotted boxes highlight regions where the MagViT-v2 fails catastrophically, like faces and hands, which have a lot of artifacts in the reconstructions. (right) We show that our highly compressed latent space (16$\times$) is suitable for training high-quality text-to-video diffusion models.
@@ 1 Introduction
The emergence of diffusion models [16, 39] has transformed image [36, 33, 10, 35, 30] and video [11, 5, 15, 38, 17, 46, 6] generation. These models gradually transform random noise into coherent visual outputs, showcasing impressive capabilities in producing high-fidelity content [6, 9]. Latent diffusion models (LDMs)[35] have gained popularity for image and video generation[36, 2]. Unlike pixel diffusion models that operate directly on raw pixels, LDMs project pixels into a low-dimensional latent space using a variational autoencoder (VAE) [24], enabling diffusion in this compact space. This dimensionality reduction enhances computational efficiency, which is particularly crucial for video generation due to the higher dimensionality of video data compared to images.
Early latent video diffusion models (LVDMs) relied on latents extracted from each frame independently [38, 15, 3] by reusing the same VAE from image LDMs. This neglects the temporal dynamics inherent in video data, compromising the temporal consistency of the generated videos, and offers no temporal compression in the latent space. To address these limitations, the seminal work MagViT-v2 [47] pioneers a spatiotemporal video tokenizer that can jointly encode images and videos in the same latent space. Originally designed to encode video into discrete tokens, MagViT-v2 has been widely adapted to continuous tokens for use in many state-of-the-art LVDMs [14, 46, 51, 26]. In the context of this work, we use MagViT-v2 as a continuous tokenizer.
@@
Interestingly, we observe that 4$\times$ temporal compression MagViT-v2 can accurately reconstruct videos with 4$\times$ frame subsampling (i.e., low FPS), much better than a 16$\times$ temporal compression model on the original (high FPS) video. Thus, we try to answer the question: “Can we meaningfully boost the temporal compression of video tokenizers by reusing a pretrained 4$\times$ model, while keeping the number of latent channels constant?” In this paper, we intentionally keep the latent channel dimension fixed over the progressive growing process for two reasons. First, it helps separate the effects of channel dimensions and progressive growing on the model performance. Second, one of our main goals is to achieve good latent features not only for high reconstruction quality but also for diffusion model training. It has been observed that increasing the latent channel dimensions, while leading to a slight improvement in reconstruction, tends to make generative model training more difficult [47].
@@
  • We present our progressive model growing technique to train ProMAG for high temporal compression ratios, achieving up to 8$\times$ and even 16$\times$. Notably, we are the first to achieve high-quality reconstruction with a 16$\times$ temporal compression model. Evaluations on video benchmarks show that ProMAG delivers significantly better reconstruction quality than simply extending existing tokenizers for higher compression. Our method also consumes 2.7$\sim$ 2.7$\times$ 2.7 × less training time compared to full model training.
  • We showcase the effectiveness of our highly compressed latents for text-to-video generation with DiT. Experiments on the text-to-video evaluation benchmark [19] reveal that DiT trained with our 16$\times$ temporal-compression latents achieves comparable or slightly higher quality than the standard 4$\times$ compression, while significantly enhancing efficiency and token length.
@@ 2 Related work
Latent video diffusion models. The pioneering LVDMs [3, 38, 15, 11, 4, 29] repurposed image VAEs to transform each frame of the video independently into the latent space. This caused temporal flickering in the generated video and limited the video length due to the limited compression achieved by image VAEs. More recent works [14, 46, 51, 26, 6] use VAEs that operate on video-level, hence mitigating both limitations of using image VAEs for video diffusion models. Building upon this, we aim to create a VAE with even higher temporal compression (8$\times$ and 16$\times$) compared to the contemporary 4$\times$, enabling highly efficient and longer video generation.
@@
High compression VAEs. There have been few works recently that investigate very high compression in latent space for higher-resolution image and longer video generation. UltraPixel [34] provides an image autoencoder that can perform 24$\times$ spatial compression. Very recently, DC-AE [8] provide a solution to increase the spatial compression to as large as 128$\times$ spatially. However, they also have to increase channel dimension consistently to preserve reconstruction quality. MovieGen [31] and Cosmos [1] also achieve 8$\times$ temporal compression by making modifications to the tokenizer architecture. In comparison, we tackle a different problem in this work. Starting from a well-trained 4$\times$ temporal compression model, our goal is to create a method to increase the temporal compression while keeping the number of latent channels fixed to avoid making the latent space harder to learn for the diffusion model. In this work, we keep the spatial compression fixed at 8$\times$. To the best of our knowledge, we are also the first method to achieve 16$\times$ temporal compression.
@@
Figure 2: Motivation. We highlight the key motivation of our progressive growing approach. Directly training MagViT-v2 for 16$\times$ temporal compression leads to poor reconstruction quality for a 242424-fps video (bottom). $s_{f}$ stands for frame subsampling factor. However, we observed that the 4$\times$ temporal compression model can still accurately reconstruct a 666-fps video by feeding the same 242424-fps video after subsampling frames by a factor of 4×4$\times$4 ×, $s_{f}=4$ (top). This observation implies that it is not necessarily the large motion that leads to worse reconstruction, but that training many downsampling (upsampling) layers of encoder (decoder) at once makes training difficult, as also observed in DC-AE [8].
Hierarchical models. Hierarchical generation has been explored most notably using generative adversarial networks (GANs) [13] to perform high-resolution image generation in stages [45, 21, 20]. These approaches first learn a generator that operates in low resolution, followed by learning additional high-resolution blocks on top in stages. Similar ideas have also been explored in image and video generation using diffusion models [34, 2, 17, 10], where content is first generated in low resolution and then upsampled in the spatial and temporal domain using a separate model. In our work, we are inspired by ProGAN [20] and adapt the progressive learning idea to our problem of boosting the temporal compression in video tokenizers, progressively learning separate model blocks to handle different levels of compression.
@@ 3 Method
Our goal is to create a continuous space video tokenizer with high temporal compression ratios (8$\times$ or 16$\times$) and good reconstruction quality. In Section 3.1, we discuss the modifications we make to the MagViT-v2 [47] to arrive at our (ProMAG) base model. In Section 3.2, we analyze why the base ProMAG, capable of doing 4$\times$ temporal compression, has difficulties extending to 8$\times$ or 16$\times$ temporal compression. We discuss a method of progressively growing the ProMAG-4$\times$ model to achieve 8$\times$ and consequently 16$\times$ temporal compression in Section 3.2. Finally, in Section 3.3 we introduce our layer-wise spatial tiling technique during decoding to enable high-resolution encoding-decoding.
@@ 3.1 Base model: ProMAG
Efficient upsampling. MagViT-v2 encodes a 17-frame video into 555 latent frames which are decoded to 202020 frames after 4$\times$ temporal upsampling. The presence of three additional frames in the reconstruction is a byproduct of the fact that the first latent frame only represents a single input frame. The MagViT-v2 decoder discards these 333 reconstruction frames when generating the final output. We find this operation to be wasteful in terms of computation and becomes a greater problem with higher temporal compression. For example, in the case of 16$\times$ compression on 171717 frames, the decoder will output 323232 frames and have to reject the first 151515 frames. To mitigate this issue, in all the temporal upsampling layers of the decoder, we discard the $1^{st}$ frame, as it can be considered a padding frame in causal convolutions. This reduces the memory consumption significantly, especially for higher compression models, without hurting the reconstruction quality.
Removing ‘spot’ artifacts. MagViT-v2 uses group normalization to normalize activations between layers. This causes ‘spot’ like artifacts when trained to progressively increase the temporal compression (Figure 3(a)). This occurs because the model tends to encode important global information in these high-norm latent pixels, as also observed in [31]. To address this phenomenon, we modify the group-normalization operation by removing the mean subtraction following [22, 23]. We found this effectively resolves the ‘spot’ artifacts in the reconstructed videos (Figure 3(b)).
@@ 3.2 Progressive model growing
Motivation. To increase the temporal compression of ProMAG from 4$\times$ to 8$\times$ or 16$\times$, the standard approach is to add additional downsampling and upsampling blocks in the bottleneck layers of the encoder and decoder respectively. In our preliminary study, we found that directly training MagViT-v2 for 16$\times$ temporal compression leads to poor reconstruction quality for a 242424-fps video (Figure 2(b)). However, we observed that the 4$\times$ temporal compression model can still accurately reconstruct a 666-fps video by feeding the same 242424-fps video after subsampling frames by a factor of 4×4$\times$4 × (Figure 2(a)). Since our base model ProMAG at 4$\times$ temporal compression is similar to MagViT-v2, this conclusion is also applicable for ProMAG.
Since the ProMAG-4$\times$ can accurately reconstruct a 666-fps video, our strategy is to use it as guidance for the additional bottleneck downsampling and upsampling blocks to achieve 8$\times$ and 16$\times$ temporal compression. This can be thought of as a guided video interpolation problem. Instead of forcing the bottleneck compression layers to learn the entire information flow, we aim to induce them to reuse the information of the 4$\times$ blocks, and only learn the essential information needed to synthesize the in-between frames, which is necessary to reconstruct the full video.
@@
Key-frame embedding. A naive way to achieve 8$\times$ temporal compression using the 4$\times$ model, $\mathbf{E}_{4\times}$, is to subsample the input video frames $\mathbf{v}$ by a factor of two, $\mathbf{v}_{//2}$, and then encode the frames with $\mathbf{E}_{4\times}$. This can be referred to as learning the key-frame encodings $z_{\text{key}}$ of the subsampled keyframes.
@@
where $\mathbf{\hat{E}}_{4\times}$ is the encoder blocks without the bottleneck $\times 1\times 1$ 1 layers in Figure 4.
Residual embedding. Now, with the encodings to accurately reconstruct keyframes in the video, we only need to learn the information for the remaining frames. We denote $\overset{*}{z}$ as the embedding of entire video $\mathbf{v}$ upto the 4$\times$ temporal compression blocks, $\mathbf{\hat{E}}_{4\times}$. To reach 8$\times$ compression, we compress $\overset{*}{z}$ through the additional 2$\times$ downsampling block. However, we need to ensure that the downsampling blocks $\mathbf{\hat{E}}_{4\times\text{-}8\times}$ are aware of the information already present in $z_{\text{key}}$, as it only needs to preserve the remaining information to avoid redundancy. For this, we use adaptive group normalization (AdaNorm) to condition the intermediate latent, $\overset{*}{z}$, on the key-frame embeddings $z_{\text{key}}$ before passing it to $\mathbf{\hat{E}}_{4\times\text{-}8\times}$.
@@
For the decoder we only add a bottleneck upsampling block.
@@
We initialize the weights for the 8$\times$ temporal compression model with the 4$\times$ models and freeze the encoder and decoder blocks initialized from the 4$\times$ model. At this stage, we only train the newly added blocks and the $\times 1\times 1$ 1 bottleneck layer using (1) the $L_{1}$ reconstruction loss $\mathcal{L}_{\text{rec}}$, and (2) the KL-divergence loss $\mathcal{L}_{\text{KL}}$. While GAN loss can slightly improve the FVD score in the reconstructed videos, it tends to introduce visual artifacts in challenging scenarios. Therefore, we do not use it at this stage, which also makes training with progressive growing significantly faster. Details about training efficiency are provided in the supplementary.
@@ 3.3 High resolution video reconstruction
Directly using our model, ProMAG, or MagViT-v2 to encode and decode videos at high resolution (e.g. 540$\times$960), results in out-of-memory (OOM) errors due to the VRAM-intensive 3D convolution layers. We found the OOM issue arises only during decoding (i.e., high-resolution videos can be encoded as-is). With this in mind, we opted to encode the entire video at once and only spatially tile the encoded latent, decoding each tile separately. However, similar to tiling in RGB space, this approach results in artifacts, as seen in Figure 5 (a). In the latent tiling case, we found the issue to be a misalignment between the full, low-resolution video encoding and decoding regimen learned during training and the tiled decoding seen at inference. Specifically, we found that there was a meaningful difference in normalization statistics encountered during training and inference. To mitigate this discrepancy, we introduce a layer-wise spatial tiling technique: during decoding, we divide the input tensor to each conv layer into overlapping spatial tiles, process each tile independently, and merge the tiles back together using linear interpolation weights to blend the overlapping components. Using this approach, we can decode into high-resolution videos without any artifacts (Figure 5 (b)).